X Company plans to prepare annual financial statements. After the company's first year of operation, the following financial statement information is available:

Balance Sheet
  Total assets $14,450
  Total liabilities 8,579
  Total Paid-In Capital 4,825
Income Statement
  Total expenses $20,335

What was total revenue for the year?
Answer #1
Total Asset =Total Liabilities +Total Paid in Capital + ( Revenue - Expense)
14450=8579+4825+(Revenue-20335)
14450=13404+ (Revenue-20335)
1046= Revenue- 20335
Revenue= 20335+1046=$21381